### v1

# 3D-ADAM: A Dataset for 3D Anomaly Detection in Advanced Manufacturing

Abstract
Surface defects are one of the largest contributors to low yield in the manufacturing sector. Accurate and reliable detection of defects during the manufacturing process is therefore of great value across the sector. State-of-the-art approaches to automated defect detection yield impressive performance on current datasets, yet still fall short in real-world manufacturing settings and developing improved methods relies on large datasets representative of real-world scenarios. Unfortunately, high-quality, high-precision RGB+3D industrial anomaly detection datasets are scarce, and typically do not reflect real-world industrial deployment scenarios. To address this, we introduce 3D-ADAM (3D Anomaly Detection in Advanced Manufacturing), the first large-scale industry-relevant dataset for high-precision 3D Anomaly Detection. 3D-ADAM comprises 14,120 high-resolution scans across 217 unique parts, captured using 4 industrial depth imaging sensors. It includes 27,346 annotated defect instances from 12 categories, covering the breadth of industrial surface defects. 3D-ADAM uniquely captures an additional 8,110 annotations of machine element features, spanning the range of relevant mechanical design form factors. Unlike existing datasets, 3D-ADAM is captured in a real industrial environment with variations in part position and orientation, camera positioning, ambient lighting conditions, as well as partial occlusions. Our evaluation of SOTA models across various RGB+3D anomaly detection tasks demonstrates the significant challenge this dataset presents to current approaches. We further validated the industrial relevance and quality of the dataset through an expert labelling survey conducted by industry partners. By providing this challenging benchmark, 3D-ADAM aims to accelerate the development of robust 3D Anomaly Detection models capable of meeting the demands of modern manufacturing environments. We provide our dataset for accessibility at: https://huggingface.co/datasets/pmchard/3D-ADAM and benchmark source code at: https://github.com/PaulMcHard/3D-ADAMBench
1 Introduction
The manufacturing sector constantly strives to reduce costs and improve quality and consistency through the use of advancements in technology. A key goal is to automatically control and monitor the entire manufacturing life-cycle of a given part. This includes detecting any anomalies or defects that may arise during the manufacturing process, and taking appropriate corrective action scheer2015 .
Accurately and reliably detecting surface defects during the manufacturing process remains a significant challenge across the manufacturing industry apqc2024 , affecting over 42% of manufacturers. 2D Anomaly Detection methods are a mature and well established technology, with a diverse range of high-volume 2D Anomaly Detection Datasets available mvtec2021 SMAP2018 and a large body of work focussed on tasks such as few shot anomaly detection WinCLIP2023 AnomalyCLIP2023 and unsupervised anomaly detection DFM2024 ContextFlow++2024 . In contrast, 3D Anomaly Detection methods lack the same degree of maturity. We attribute this in part to the fact that suitable datasets are limited and do not provide the breadth of features and defect examples required for developing solutions capable of performing at an acceptable level in advanced manufacturing contexts. Three leading datasets have been proposed previously for the challenge of 3D Anomaly Detection, MVTec3D-AD mvtec2022 , Eyecandies eyecandies2022 and Real3D-AD real3dad2023 . Each of these datasets has several limitations which make them unsuitable for industrial applications, with the common issue between them being a lack of breadth in the defect and feature categories presented in manufacturing samples.
Hence, we introduce 3D-ADAM, a large-scale, high-resolution, multi-camera dataset for 3D-Anomaly Detection in Advanced Manufacturing. 3D-ADAM addresses the limitations of existing datasets by including comprehensive suite of examples of defects in 12 classes across a range of machine element features Mott2023 , which cover the range of surface defect types that may be encountered in precision part manufacturing environments. 3D-ADAM proposes a set of 16 machine elements Mott2023 features across all part categories, with care taken in curating a range of defect examples displayed on each machine element type.
The manufactured parts contained in 3D-ADAM constitute a complete assembly for a robot arm mechanism, based on the open-source BCN3D-Moveo system BCNMoveo2022 and augmented with additional components sourced from the Thingi10K Thingi10K2016 part dataset in order to ensure all relevant machine elements are adequately represented. We manufactured all the parts, both good and defective, in Polylactic Acid (PLA) as it is amongst the most commonly used materials for Fused Deposition Modelling (FDM) based additive manufacturing. Some defects naturally arose during the printing process, while others were forced by adjusting the printing parameters or by manual intervention. We assess the performance of our dataset in the tasks of 3D and RGB + 3D Anomaly Detection in a supervised setting across a suite of leading models, demonstrating that the 3D-ADAM dataset presents a novel challenge for current leading architectures in 3D and RGB+3D Anomaly Detection, which will bridge the gap between the current SOTA in 3D Anomaly Detection models and the level of performance required for deployment in real-world scenarios. Our main contributions are:
  • We propose the first multi-camera, multi-modal, industry-relevant dataset for 3D and RGB+3D Anomaly Detection and Localization. Comprising 14,120 high-resolution image and 3D point cloud scans across 28 categories, with 27,346 annotated defect regions in 12 classes and a further 27,346 annotated machine element features in 16 classes (Figure 1).
  • We evaluate leading methods in supervised 3D and RGB+3D Anomaly Detection and Localisation. Our initial benchmark demonstrates that these methods under-perform on our dataset, and that there is significant room for improvement for such methods to be suitable for real-world environments.
  • We evaluate the quality of our defect segmentation mask annotations through a survey of industry experts which shows that individuals with expert knowledge of surface defects in advanced manufacturing contexts strongly agree with the image-level quality and pixel-level accuracy of the annotations proposed in our dataset.
Figure 1: Highlighted ground truth defect examples across each defect class.
2 Related Work
3D Anomaly Detection Datasets.
Three datasets have previously been proposed for 3D Anomaly Detection, MVTec3D-AD mvtec2022 , Eyecandies eyecandies2022 and Real3D-AD real3dad2023 . Of these, MVTec 3D-AD and Real3D-AD represent the current primary datasets used as benchmarks for the problem of 3D Anomaly Detection in manufacturing. The Eyecandies dataset eyecandies2022 , while a novel proposal in its own right, comprises a set of synthetically generated parts and defects on object instances that are unsuitable for industrial contexts, and is therefore not considered a viable dataset for such real-world applications. In their proposal for the Real 3D-AD Dataset, Liu et al. real3dad2023 claim their dataset is more suitable for achieving high-precision 3D Anomaly Detection. Indeed, this claim is supported by their notably higher point resolution and precision when compared to the MVTec 3D-AD dataset, which is a result of the difference in point-cloud acquisition hardware used. However, the MVTec dataset remains far more popular among new techniques published for 3D Anomaly Detection. Both datasets feature a similar number of object categories, 10 in the MVTec 3D-AD Dataset and 11 in the Real3D-AD Dataset, respectively, with the volume of pointcloud scan data per object category also comparable. The key distinction between the two is that the MVTec dataset presents a superior and far more rigorous ground truth annotation methodology and ontology, and for this reason remains the more widely adopted.
Industrial Anomaly Detection.
Advances in computer vision and deep learning are leading the technological drive in improvements to defect detection in the manufacturing industry henriksson2020 . Saeed et al saeed2021 , for example, demonstrate that frameworks such as Faster Regional-Convolutional Neural Networks ren2017 , Faster R-CNN, perform strongly in solving the task of fault detection and small object detection in an industrial environment. The foundational work in Deep Learning on 3D Point Clouds is the PointNet qi2017 architecture, which proposed a novel architecture for classification and segmentation tasks, taking sets of unordered 3D points from point clouds as input. Since PointNet was proposed, it has provided the basis from which many deep learning challenges have been handled in 3D point cloud data. Its successor network, PointNet++ qi2017plus , introduced the idea of applying PointNet recursively to nested partitions of the input point set. The introduction of this approach allows the network to learn local features of the part by exploiting space metrics to develop an understanding of the context surrounding a region. Not only does this underpin some of the state-of-the-art 3D Anomaly detection models shapeGuided2023 , but it has also served as the foundation for a number of novel techniques and architectures that look to expand beyond it.
The current state-of-the-art in 3D Anomaly Detection cpmf2024 3dsr2024 shapeGuided2023 transfusion2024 presents a number of highly capable models for unsupervised 3D anomaly detection. These models have all been trained, tested and validated on the MVTec 3D-AD Dataset mvtec2022 . Each of these models is underpinned by backbones built on the PointNet qi2017 and ResNet resnet2015 models, however vary significantly in the architectures proposed. Chu et al. shapeGuided2023 propose a model for Shape-Guided expert learning, in which two separate architectures are trained for 3D feature extraction, based on PointNet part segmentation, and 2D anomaly detection, based on ResNet-based resnet2015 RGB feature mapping, respectively. The 3D features then act as a guide for the RGB features to build dual memory banks from which inference can be conducted for anomaly detection in an unseen part. Zvartanik et al. 3dsr2024 propose the Depth-Aware Discrete Auto-Encoder (DADA) architecture, which takes both the RGB and Depth image data jointly as a single input, enabling the model to learn in a general discrete latent space. The DADA architecture is the core of their 3D Dual Subspace Reprojection Network (3DSR) 3dsr2024 , which is an extension of the authors’ previous work zavrtanik2022 , in using dual-encoder models for 2D anomaly detection. Conversely, a leading model in 3D Anomaly Detection proposed by Cao et al. cpmf2024 does not rely on raw RGB information as input in any capacity. Instead, they propose a Pseudo Multi-Modal Feature, in which 2D features are extracted by creating projections of the 3D point cloud data into 2D at a given number of view-points, and relying on pre-trained 2D image networks. Their model, based on ResNet, is similar to Chu et al.’s shapeGuided2023 Shape-Guided architecture. These multi-view features are then aggregated together with hand-crafted 3D features, to construct the memory banks from which inference can be performed for anomaly detection. Lastly, Fucka et al. transfusion2024 propose a model that leverages modern diffusion-based techniques in a transparency-based diffusion model for 3D Anomaly Detection, with promising results.
On the MVTec 3D-AD Dataset upon which the leading models are all trained, Fucka et al.’s TransFusion model transfusion2024 , currently has the hightest performing results on both Detection and Segmentation tasks in RGB+3D Anomaly Detection, whilst Chu et al.’s shapeGuided2023 Shape-Guided model performs best on Segmentation benchmarks, and Cao et al.’s 3dsr2024 CPMF model provides the strongest performance on Detection Benchmarks, amongst available models. It should be noted that this description excludes the Global Local Feature Matching (GLFM) model proposed by Cheng et al. cheng2025 , as while this model outperforms both Shape-Guided and CPMF at the time of writing, it is not yet publicly available.
3 The 3D-ADAM Dataset
The 3D-ADAM dataset consists of 12,192 scans distributed equally between 4 high-resolution industrial 3D sensors. The sensors deployed are MechMind LSR-L, MechMind Nano, Stereolabs Zed 2i, and Intel Realsense (as shown in Figure 2). The details of these sensors are described in Table 1. To obtain a comprehensive coverage of the technologies deployed in industrial vision applications, these sensors vary in quality, cost, capture resolution, and depth-sensing technology.
Sensor LSR-L Nano RealSense D455 Zed2i
Manufacturer Mech-Mind Mech-Mind Intel Stereolabs
Sensor Type Sony CMOS Sony CMOS, Intel SDM Stereo Depth Sensor
Resolution (pix) 2048 x 1536 1280 x 1024 1280 x 720 1920 x 1080
Optimal Range (mm) 1500 - 3000 300 - 600 600 - 6000 300 - 1200
FOV (mm) 1500 x 1200 @ 1.5m 220 x 150 @ 0.3m 825 x 580 @ 0.6m 490 x 345 @ 0.3m
3000 x2400 @ 3.0m 440 × 300 @ 0.6 m 8260 x 5820 @ 6m 1965 x 1375 @ 1.2m
Precision (mm) 0.5-1.0 0.1 80 30
Table 1: Specifications of sensors used to capture the 3D-ADAM dataset.
We aim to provide exhaustive coverage of the range of relevant defect and machine element types, such that we offer a diverse range of defect examples for industrial anomaly detection applications. To achieve this, 3D-ADAM contains scans of parts from 29 distinct object categories. The majority of these parts represent the complete assembly for a robot arm mechanism, obtained from the open-source BCN3D-Moveo system BCNMoveo2022 . This set is augmented with additional components sourced from the open-source Thingi10k dataset Thingi10K2016 to provide a full suite of machine elements, covering common elements such as faces and edges, internal and external fillets, internal and external chamfers, holes, kerfs, tapers, indents, counterbores, countersinks as well as the most common gear types; spur gear teeth, rack gear teeth, spiral gear teeth and both clockwise and counter-clockwise helical gear teeth, for a total of 16 distinct machine element classes (as shown in Figure 1 in the supplementary material). In addition, we curate a full suite of defect examples across this breadth of machine element examples, distributed across the range of part categories. The range of defect classes is designed to represent the full spectrum of distinct surface defect types which can be encountered in manufacturing environments and includes (shown in Figure 1); cuts, bulges, holes, gaps, burrs, cracks, scratches, marks, warping, roughness and, specific to the additive manufacturing process employed in producing this dataset, over extrusion and under extrusion defects, for a total of 12 defect classes, which is the largest range provided by any dataset to date. Table 2 describes the breakdown of total parts, total scans and total number of defect and machine element annotations, as well as the number of unique defect classes and machine element classes featured in each category of our dataset.
Category Parts Images Defects Machine Elements # of Defect Classes # of Machine Element Classes
Robot Base Joint 10 672 600 600 2 6
Robot Base Body 11 704 1215 1215 6 8
Robot Base Motor Housing 16 972 2550 2550 8 6
Robot Shoulder Body 6 384 718 718 4 6
Robot Shoulder Joint (Left) 7 452 1404 1404 5 6
Robot Shoulder Joint (Right) 7 448 1336 1336 4 6
Robot Forearm Body 5 448 2152 2152 6 4
Robot Elbow Joint (Left) 6 384 1660 1660 4 5
Robot Elbow Joint (Right) 9 580 2614 2614 6 8
Robot Wrist Body 8 652 2085 2085 6 7
Robot Wrist Joint (Left) 9 576 1920 1920 5 8
Robot Wrist Joint (Right) 8 512 939 939 6 7
Robot Shoulder Fixture (Left) 6 384 54 54 2 1
Robot Shoulder Fixture (Right) 6 384 62 62 2 1
Robot Forearm Fixture 4 256 0 0 0 0
Robot Wrist Fixture 4 256 0 0 0 0
Robot Base Cover 10 608 991 991 6 4
Robot Shoulder Cover 12 688 820 820 5 5
Robot Elbow Cover 8 448 527 527 2 5
Robot Wrist Cover 8 480 162 162 2 4
Base Clamp Body 2 124 0 0 0 0
Base Clamp Bolt 2 132 0 0 0 0
Robot Gripper 11 896 338 338 4 3
Helical Gear (CW) 6 384 434 434 2 3
Helical Gear (CCW) 9 548 731 731 4 3
Rack Gear 8 516 1146 1146 7 4
Spiral Gear 9 560 1314 1314 5 5
Spur Gear 10 672 1574 1574 6 4
Total 217 14120 27346 27346
Table 2: Statistical Breakdown. # of defect and # of machine element classes is given per per part category.
3.1 Structure & Data Acquisition
We employed additive manufacturing via Fused Deposition Modelling (FDM) cano2021 to manufacture the parts included in our dataset. For this task, four distinct systems were employed: two Prusa i3 systems, an Ender 3 Pro, and an Ultimaker 2, and all parts were manufactured in Polylactic Acid (PLA). In the majority of instances, these systems were used to produce one complete set of "good" parts, manufactured at optimal settings. Many of these "good" examples were ultimately found to exhibit some minor degree of surface defects, discovered during the annotation stage. The set of parts for defective cases was distributed randomly and evenly amongst the four systems. Defect instances were created either by manipulating manufacturing settings on the systems, in order to create manufacturing process-based defects such as bulges, gaps, warping, over extrusion and under extrusion. Surface defect examples of a mechanical nature, in contrast to those generated by process error, were generated by replicating the handling and manipulation errors. These examples result in a real manufacturing environment, including the generation of cuts, holes, cracks, marks, and scratches. The majority of instances of burrs and roughness defects were as a result of naturally occurring defective features on parts.
Figure 2: Experimental setup in laboratory conditions, showing mouting of the MechMind Nano, Intel RealSense and Stereolabs Zed2i sensors on an ABB IRB 120 robot arm, oriented towards the mounting fixture. Not pictured: MechMind LSR-L camera, mounted out of view above the cell frame.
Data was acquired in a real-world environment in the robotics applications development cell at HAL Robotics. (Figure 2). Within this cell, an ABB IRB 120 robot arm was used to mount and orient the MechMind-Nano, Intel RealSense and Stereolabs Zed2i sensors, while MechMind LSR-L sensor was fixed to the upper exterior of the cell enclosure. The positions of all four industrial 3D imaging sensors were static throughout the data acquisition process. This mounting configuration was used to ensure that all four sensors would be operating within their optimal scanning ranges, described in Table 1, oriented towards a mounting fixture that held the parts in place during scanning. The mounting fixture allowed 2-axis rotation, with a 360-degree rotation around its primary axis and a 90-degree range of motion in its secondary axis. In each scanning procedure, the part was fixed at the centre of this fixture, and allowed to rotate through a range of orientations, with a minimum of a full rotation of the primary axis at 45-degree intervals, with the secondary axis fixed at both 0 and 30 degrees, respectively. This provided, in most instances, a complete view of the part. In those instances where additional coverage was desired, further variance in the secondary axis position was used to achieve the complete view.
Each of the industrial scanners deployed captured six-channel images in the capture resolutions defined in Table 1, with these channels representing the r, g, b, x, y and z coordinates with respect to the sensors’ local reference coordinate frame. These [x, y, z] vectors represent the point cloud scan obtained by the sensor, and in all instances are captured as .PLY files. Additionally, for all sensors, the corresponding [r, g, b] values for each pixel are captured as a PNG RGB image, which enables us to have a 1-to-1 mapping between the RGB and XYZ image data for each scan. The scene captured within our environment was open to indirect natural lighting sources during the period of scanning, as well as diffuse indirect artificial light sources when natural illumination levels were insufficient for proper data acquisition conditions. Thus, as scanning was conducted throughout the day, across weather and seasonal change variations, the data captured represents a wide range of natural and artificial lighting conditions as can be observed in Figure 1 and 3.
Figure 3: Grid (right): Sample of images shown for selected parts in the subset shown, demonstrating differences in camera resolution, orientation and variations in lighting conditions. Images in row A were captured using the MechMind LSRL, while B, C and D, using the MechMind Nano, Intel RealSense and Stereolabs Zed 2i, respectively. Left: A sample subset of the dataset arrayed for capture on a table adjacent to the robot cell.
3.2 Ground Truth Annotations
We provide accurate, high-detail ground truth annotations for each instance of a part, machine element and defect in dataset. Part annotations are provided as a segmentation mask, machine element annotations are a bounding box with a class label and defect annotations are a segmentation mask with a class label. These are demonstrated for an example part in Figure 4.
For all cameras, a semi-automated segmentation method, based on Cutie cutie2023 , was employed to create foreground masks isolating the part in the scene. These were checked by hand and corrected where necessary. Defect and Machine Element ground truth annotations required expert knowledge of the manufacturing defects present in a part, and the machine elements on which they occurred. For this reason, all defect segmentation masks and machine element bounding boxes were annotated by hand on the 2D image data of each scan using the set captured from the MechMind Nano, as this sensor provided the highest level of detail of defects. A single expert annotator performed the annotations for all images in the MechMind Nano set, with verification and feedback support provided by HAL Robotics experts. Homographic transform methods were employed to complete the ground truth annotations. These methods calculated the transformation matrix between part masks from the Mech-Mind Nano image to the equivalent from each of the remaining sensors, applying this transform to the defect masks and machine element bounding boxes to propagate the annotations from the MechMind Nano image set to the remaining systems with these being checked by hand with minor corrections performed where necessary.
Figure 4: Visualization of Part Segmentation Mask, Defect Segmentation Masks and Machine Element Bounding Box annotations.
4 Benchmark and Evaluation Methodology
We evaluate our dataset against a suite of models in a supervised learning setting at the tasks of 3D and RGB+3D Anomaly Detection and Localisation. For this task, we divide our dataset into 3 sets, an anomaly-free training set, containing all of the defect-free scans captured for the dataset. The annotated defective scans are partitioned into a test and validation set in a 60:40 share. To evaluate how existing 3D Anomaly Detection and Localisation models perform on our dataset, as well as assess the challenge presented by the dataset, we carried out a benchmark on the current state-of-the-art models. For this, we include the four currently leading or recently leading available models at 3D and RGB+3D Anomaly Detection and Segmentation task on the MVTec3D-AD Dataset mvtec2022 paperswithcodergb3d paperswithcode3d . These are Transfusion transfusion2024 , CPMF cpmf2024 , 3DSR 3dsr2024 and Shape-Guided shapeGuided2023 .
Evaluation Methodology.
To provide a strong foundation for comparing the performance of the models included within the benchmark set on our dataset at the tasks of 3D Anomaly Detection and localisation, standard evaluation metrics for anomaly detection are employed. This maintains a consistent evaluation criteria across models, which allows accurate comparison of each model’s performance with respect to the MVTec 3D-AD Dataset, as well as comparison of the performance of each respective model on our dataset. These standard criteria are defined on the image-level anomaly detection performance and the pixel-level anomaly localisation performance. For image-level performance, the Area Under Receiver Operator Curve (AUROC) metric is utilised, while for the pixel-level anomaly localisation the Area Under Per Region Overlap (AUPRO) metric is used.
5 Industry Expert Survey
To validate the accuracy and suitability of our annotation methodology for industrial applications, we conducted a survey with industry experts, who were asked to re-annotate a subset of our data. Five industry experts from three different manufacturing organisations, across the industrial research and development, digital process manufacturing and robotics sectors from both private and public sector organisations, volunteered to take part in this survey. Each was provided with a randomised set of defective scans representing 1% of the total dataset, with the intention that the task should be completable by a single expert in a few hours. A detailed set of instructions for the labelling task was provided, also included in our github repository, as well as "good" examples for the images assigned, providing as-designed, defect-free examples of each part instance alongside the defective case. Other than those images included in the instruction guide, participants were not provided with any examples of annotated images from the core dataset, in order to minimise the degree of influence on their approach to annotations, and provide as close to a blind evaluation of the annotation process as possible. The ground truth annotations produced by experts were then evaluated for similarity to the equivalent ground truth annotations from the core dataset using a similar methodology as Section 4. However, given all labelling is conducted based on 2D image data, in place of the AUROC and AUPRO methods employed for Image-Level and Pixel-Level comparison respectively which is appropriate in 3D contexts, we instead use the Intersection over Union (IoU) criteria for image-level comparison and the PRO (Per Region Overlap) criteria, which is the core of the AUPRO method and more appropriate for 2D contexts.
Figure 5: Comparison of Pixel-Level annotations between dataset ground truth and labeller annotations. Light Blue: True Positive; Red: False Negative; and Green: False Postive.
The results of the labelling survey demonstrate strong correlation between the annotations proposed within our dataset and those produced by industry expert surveys. Mean IoU score across all surveyed experts was 0.6 with a standard deviation of 0.04, indicating good image-level agreement on overall defect annotations for a given sample, with scores amongst labellers being tightly clustered. A mean PRO score of 0.76 with a standard deviation of 0.058 indicates strong pixel-level agreement on the shape and location of individual defects, with results again exhibiting low variability between labellers. Figure 5 provides a visual example where we have taken an instance where the task sets of two labellers overlapped for this single example. Alignment between labellers and the ground truth is strong, with labellers demonstrating strong correlation of true positive pixels, and tending to favour false positives over false negatives, thereby indicating that industry experts scrutinised a slightly more rigorous threshold for considering a given region to be defective in comparison to that proposed in our dataset.
6 Discussion
The results of the benchmarking evaluation and are included in the supplementary material. We find that across the set of models tested with our dataset, the results indicate that the 3D-ADAM dataset provides a significantly greater challenge in comparison to the performance of the same models on the MVTec3D-AD dataset. The 3D-ADAM dataset presents additional new challenges in 3D Anomaly Detection, Localisation and Classification which are not currently facilitated by available datasets. Such challenges include few-shot and zero-shot multi-modal anomaly detection and localisation, multi-modal anomaly classification, and machine element feature classification. Additionally, Unsupervised Anomaly Detection tasks can be performed with our dataset, though this is not explored here and presents an exciting opportunity for future work. No current datasets exist which are appropriate for unsupervised settings in 3D and RGB+3D Anomaly Detection.
Contribution.
The 3D-ADAM dataset is demonstrably more challenging than those preceding it. It proposes a uniquely comprehensive range of both defect types, covering a breadth of industrial defect types with instances presented across a range of annotated machine elements, which is a crucial factor for manufacturing applications and something no prior dataset has attempted. This dataset also uniquely proposes images and 3D scans of machine parts captured in-the-wild from multiple distinct 3D imaging sensors, further exemplifying its suitability for developing solutions suitable for real-world industrial deployment in advanced manufacturing use cases.
Broader Impact and Ethical Concerns.
Given the importance of identifying defects in manufacturing environments as a sector-wide challenge, the development of novel solutions for anomaly detection based on 3D image data is crucial for delivering a new generation of technologies for fully automating the detection, classification, localisation and ultimately correction of manufactured parts, in-process. The results of our benchmarking evaluation demonstrate that current leading models are not well equipped for this task due to the limitations of existing datasets for training, testing and validation. Our model provides a significant increase in the challenge presented, bringing it in line with realistic real-world use-cases that would be faced on deployment, thereby contributing to the broader ambition of developing mature 3D anomaly detection solutions in line with what is currently available in 2D anomaly detection. We anticipate no concerns around security, personal data protection, violation of intellectual property rights or any other major risks with this work. All part models included are obtained from open-source and freely licensed repositories, the additive manufacturing process was conducted using commercial-grade systems and materials and scanning was conducted in a closed cell without people present. The industry expert survey was conducted anonymously and no details pertaining to the experts involved or their individual contributions we retained once the results of the survey had been obtained.
Limitations.
We acknowledge that opportunities for improvement exist beyond what is proposed by our dataset. By composing our dataset exclusively of PLA-based additively manufactured parts, we present only one type of texture for models to identify surface defects upon. In many high-value advanced manufacturing applications, subtle differences in the presentation of certain defects may occur due to differences in material properties. The degree this would impact applications in real-world use cases has not been explored and presents an interesting opportunity for future study. Additionally, while the range of surface defects accounted for in our dataset is the most comprehensive of any dataset for 3D Anomaly Detection to date, there remains an array of potential defects which can occur in advanced manufacturing use-cases, well understood by experts in the manufacturing and materials science fields, which are not provided in this dataset. This presents an opportunity for further augmentations to accommodate a broader range of industrial applications. Lastly, our dataset does not provide full 3D views of the parts included, as some datasets which precede it do, however we do not consider this a significant limitation, as we consider the availability of such complete scans to be an unrealistic standard to expect for real-world applications and instead anticipate that the multi-view, multi-camera in-the-wild approach proposed is significantly more appropriate for such instances.
7 Conclusion
In this work, we present 3D-ADAM, a novel dataset for 3D anomaly detection, which aims to present a challenge closely aligned with the scenarios encountered in real-world industrial applications for Advanced Manufacturing. 3D-ADAM is the largest dataset for this task to date, hosting the broadest range of defects of any dataset for 3D anomaly detection. 3D-ADAM uniquely provides machine element annotations for each defect, which no previous dataset has proposed, and is unique in presenting this challenge through a multi-view, multi-camera, in-the-wild dataset. 3D-ADAM provides a significantly more challenging dataset than any prior dataset, and is unique in having its defect annotation methodology verified through a survey of industry experts.
Acknowledgements and Disclosure of Funding
This research was enabled in part through the manufacturing and laboratory resources, as well as the input of industry expertise provided by HAL Robotics. We also thank Jose Hernandez of the National Skills Academy, and Michael Canavan and Lewes Glen of the Industrial Design team at the National Manufacturing Institute Scotland, Andrew Hamilton of the Digital Process Manufacturing Centre, and Thibault Schwartz of HAL Robotics for their contributions in our survey of industry experts. Paul McHard has received funding from HAL Robotics and the Royal Commission for the Exhibition of 1851 via an Industrial Fellowship.